# Specialized Foundation Models Struggle to Beat Supervised Baselines

### v1

@@ 4.3 Time series
Auto-ARIMA 10 0 0.896 12.82 0.00 0.00
AR 513 0 0.556 6.57 31.17 24.31
@@
Table 3: Aggregate performance on time series tasks across seven tasks. The latter evaluation demonstrates that simply tuning a classical AR model is competitive with state-of-the-art FMs while using no pretraining data and tens of thousands of times fewer parameters. For Mean/Median %Imp. we report percentage improvement over Auto-ARIMA, and for Auto-AR, the model size refers to the largest configuration across tasks. “-” indicates unknown quantities.
Our last domain is time series, which has many FMs, including those that use the standard pretrain-then-fine-tune workflow: GPT4TS (OFA) (Zhou et al., 2023a), LLM4TS (Chang et al., 2023), MOMENT (Goswami et al., 2024), TEST (Sun et al., 2023), S2IP-LLM (Pan et al., 2024), CALF (Liu et al., 2024), TTM (Ekambaram et al., 2024), and Time-LLM (Jin et al., 2024); and others that evaluate in a zero-shot (ZS) regime: TEMPO (Cao et al., 2024), TimesFM (Das et al., 2024), Moirai (Woo et al., 2024), and Toto (Cohen et al., 2024). As we are comparing to supervised baselines, our evaluation of ZS models will be in a less challenging setting than the one they report numbers for. We study the performance of these FMs and our baselines on the problem of long-horizon forecasting, which has a standard set of tasks (Goswami et al., 2024, Table 11), of which we consider seven. [^6] The two we do not consider, Exchange and ILI, are not evaluated on by most time series FMs. Note that each task consists of four settings corresponding to different time horizons, so in total this yields twenty-eight tasks. Lastly, we compute aggregate metrics using RMSE, not MSE, so that performance scales linearly with prediction error; this choice has no effect on average rank.
@@
Note that we do not include four of the above time series FMs in our main analysis. One of them, Time-LLM, reports strong results but has had difficulty being reproduced by both us and past efforts, as we detail in Appendix C.3.1. The three others—Moirai, LLM4TS, and Toto—either evaluate on only a subset of the seven tasks or are closed-source (or both); we list their reported numbers in Table 14. As discussed in Appendix C.3.2, Moirai underperforms Auto-AR and LLM4TS performs roughly on par with TTM (A). The most recently released FM, Toto, does have strong aggregate metrics, but this is mainly due to a dominant performance on a sinle task, ETTh2 (c.f. Figure 4). As a result we do not view these excluded FMs as significantly affecting our conclusions.
4.3.1 Baselines
@@
To baseline these FMs we use mainly linear forecasting methods, including the classical (untuned) linear auto-regression (AR), the automated workhorse Auto-ARIMA (Hyndman & Khandakar, 2008), the more recent DLinear (Zeng et al., 2023), and our own workflow Auto-AR described in Section 3.2. Lastly, we also evaluate our other approach, DASHA, on six of the tasks (c.f. Table 14).
4.3.2 Results
@@
Table 3 shows that on the full seven-dataset evaluation our Auto-AR workflow always attains competitive performances across all aggregate metrics considered, and in particular attains the best median improvement over Auto-ARIMA. Specifically, our Auto-AR workflow achieves competitive performances with two other recent time series FMs, namely MOMENT and S2IP-LLM, and outperforms the rest of the FMs. Although TTM surpasses all other methods across three aggregated metrics, the improvements remain relatively marginal. This observation aligns with our assertion that the substantial increase in pretraining dataset size and model scale has not yet resulted in significant advancements in model performance. Notably, the three best performing methods are not zero-shot, which is perhaps not surprising given the extra data. However, it does reinforce the intuition that settings with high data availability should prefer supervised methods, including simple ones like AR. Notably, even our untuned implementation of AR that uses no differencing and a large lookback window is quite effective, doing better than ZS FMs across all aggregate metrics and even MOMENT on some of them.
@@ 5.2 Computational efficiency considerations
While not our main focus, we nevertheless highlight that any performance gains from FMs must be balanced against their additional cost. In addition to the extensive GPU-hours used for pretraining, the resulting models are often much bigger and so lead to much more costly inference. Indeed, apart from the special case of HyenaDNA, the CNN architectures discovered and trained using our DASHA workflow are typically over ten times smaller than FMs in the case of genomics and three to ten times smaller in the case of satellite imaging. Moreover, for time series our Auto-AR approach is quick-to-train and yields simple models with less than 1K parameters—over two-thousand times smaller than any FM—while attaining performance that is often competitive even with closed-source models. In aggregate, these examples further demonstrate the efficiency of supervised approaches and the resulting high performance bar that FMs need to clear before they can be deemed useful.
@@ 5.3 The power of tuning kernel sizes and dilation rates
Figure 3: PCA visualization of the architectures discovered for three different tasks when the DASHA workflow is run multiples times. The spatial clustering across tasks demonstrates the within-task consistency of the workflow’s architecture search component and the utility of using diverse models as baselines.
Our results for genomics and satellite imaging are driven by the DASHA workflow, whose crucial component is the tuning of kernel sizes and dilation rate in CNN backbones such as wide ResNet. Its success demonstrates that the procedure is an effective surrogate for human-driven model development, enabling the automated discovery of the types of diverse, domain-specific baselines stressed in Section 5.1. To understand this further, we study whether the architecture search component selects different kernel sizes and dilation rates for different tasks, and whether it does so in a consistent manner. Specifically, we run DASHA on three of the smaller datasets in the NT benchmark with fifteen different random seeds, construct eighteen-dimensional vectors of the discovered kernel sizes and dilation rates assigned to each of the nine layers, and project these to two dimensions using principal component analysis (PCA). The result in Figure 3 reveals that the architectures are clustered by task, demonstrating that the procedure selects different but consistent-within-task kernel parameters. This visualization suggests that architecture search is a useful surrogate for model development, and consequently that the DASHA workflow may also be useful for automating similar studies and baselining FMs in other domains with high-dimensional, unstructured data.
@@ 5.5 Limitations
While our findings are significant according to measures set by past work, they should not be misinterpreted to address all possible scenarios where FMs may be useful. This is most salient for time series FMs motivated by zero-shot concerns, a setting we do not study, and to some extent for genomics FMs, which are often used for exploratory science and not supervised learning. We are also of course computationally limited and there are many other domains where FMs have been pretrained, and even in our three there are other tasks beyond classification and forecasting. Nevertheless, our evaluation is extensive—over twenty-five FMs and over fifty tasks—and so are at least strongly suggestive of the state of a field that uses benchmark performance to motivate and justify pretraining.
@@ Acknowledgments
We thank Mononito Goswami, Esther Rolf, and Stephan Xie for useful feedback. This work was supported in part by the National Science Foundation grants IIS1705121, IIS1838017, IIS2046613, IIS2112471, a TCS Presidential Fellowship, and funding from Meta, Morgan Stanley, Amazon, Google, and Scribe. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of any of these funding agencies.
@@ C.1 Genomics
We include all FMs listed in Dalla-Torre et al. (2023, Supplementary Table 6) with addition of the two recently released models, Caduceus (Schiff et al., 2024) and Gena-LM (Fishman et al., 2024); note that the last row of tasks in the NT paper(promoter and splice sites) are mislabeled, but we infer an order in combination with previous information obtained from a (now-deleted) Huggingface leaderboard. [^7] https://huggingface.co/spaces/InstaDeepAI/nucleotide_transformer_benchmark In alignment with the leaderboard, we apply a 0.1 validation split for DASHA during our evaluation. Additionally, we use an architecture set that includes both Wide ResNet and UNet for the search with DASHA on these datasets. We use batch size$=$ 128 for all datasets, and cross entropy loss for all the training and finetuning. Individual scores for each task in the benchmark are provided in Tables 9 and 8.
@@
Table 8: Regulatory Elements and RNA Production Downstream Tasks
@@
Table 9: Histone Modification Downstream Tasks
@@
Table 11: Regulatory Elements and RNA Production for GENA-LM. (all scores are obtained using MCC) The last column shows the average MCC across all 18 datasets.
@@ C.2 Satellite imaging
Training on satellite datasets requires relatively large computational resources due to the high number of channels and the size of the datasets. To ensure a fair comparison, we fine-tuned all the foundation models ourselves by sweeping across a fixed set of base learning rates $[5e-3,2e-3,2e-4,4e-5]$. We then calculate the actual learning rate from base learning rate following previous work by $lr=base\_lr\cdot\frac{batch\_size}{256}$. This approach ensures that approximately the same amount of resources were used as during the DASHA tuning process, allowing for a balanced evaluation of model performance.
@@
It is also important to note that SatMAE only accepts 3-channel and 12-channel inputs, while CROMA is limited to 12-channel inputs. GeoBench, however, includes a wide range of tasks with varying numbers of input channels, ranging from 3 to 18. Despite these differences, we include all datasets in our evaluation because they are valuable benchmarks in the satellite image domain, and it is crucial for FMs in this field to generalize across diverse datasets. For datasets where the input size does not match the model requirements, we pad missing channels with zeros and prune any extra channels. However, to ensure a fair comparison, in addition to reporting the average scores across all datasets, we also provide average scores excluding m-pv4ger and m-forestnet, where missing channels may affect the performance of the FMs. The aggregate scores excluding m-pv4ger and m-forestnet are presented in Figure 12.
Model Average Average Mean Median
Score $\uparrow$ Rank $\downarrow$ %Imp.$\uparrow$ %Imp.$\uparrow$
Foundation Model
SatMAE-Base 77.71 6.00 6.74 4.56
SatMAE-Large 78.44 4.07 7.87 6.75
GFM 76.95 5.57 5.40 4.08
SwinT-Base 76.12 6.50 3.98 1.43
CROMA-Base 77.98 3.57 6.95 5.30
CROMA-Large 79.06 2.14 8.84 6.26
Supervised Methods
ResNet50 73.25 9.00 -0.27 -0.38
Wide ResNet 73.90 8.00 0.00 0.00
UNet 75.29 6.57 2.33 0.87
DASHA 78.00 3.57 7.02 5.16
For training and finetuning, we universally use batch_size = 16 and loss function as cross entropy with 0.1 label smoothing for single label classification, and multi-label soft margin loss for multilabel classification. Individual scores for each task are provided in Table 13.
@@ C.3 Time series
The long horizon forecasting task for a time series can be summarized as follows: at every timestep $t$, take the $L$ historical observations at times $(t-L+1,...,t)$ for each channel and predict the next $H$ observations $(t+1,...,t+H)$ for every channel. Following the literature, we evaluate models on $H\in\{24,36,48,60\}$ for the Illness dataset and $H\in\{96,192,336,720\}$. For most methods, we report results when $L=512$.
@@
TEMPO (Zero Shot) 0.211 0.254 0.292 0.37 0.178 0.198 0.209 0.279 3.0 2.956 2.651 2.701 0.476 0.496 0.503 0.538
TimesFM (Zero Shot) 0.122 0.169 0.242 0.391 0.119 0.137 0.158 0.206 2.595 2.984 3.34 3.227 0.327 0.353 0.378 0.42
@@
Table 14: Time Series Forecasting Tasks
In addition to the results for DASHA, Auto-AR, DLinear, and FMs, we evaluate the performance of two simple baselines
  1. 1.
    Vanilla Autoregressive Model (Box & Jenkins, 1976): This model predicts the (scalar) value of a time series at $t+1$ as a linear combination of the last $L$ timesteps and a constant, i.e. $\hat{x}_{t+1}=\alpha_{0}+\alpha_{1}x_{t}+\alpha_{2}x_{t-1}+...+\alpha_{L}x_{t-L+1}$ for learnable parameters $\alpha_{0},...,\alpha_{L}$. We fit these parameters using standard maximum likelihood techniques.
  2. 2.
    ARIMA is a statistical method used for time series forecasting that combines three components: AutoRegressive (AR), Integrated (I), and Moving Average (MA). The AR component models the relationship between an observation and its lagged (past) values, assuming that past values have a linear influence on future ones. The Integrated component applies differencing to the data to remove trends or seasonality, making the time series stationary by stabilizing its mean over time. The MA component models the relationship between an observation and the residual errors from a moving average model applied to previous observations. ARIMA is characterized by three parameters: p (the number of lag observations), d (the number of differencing steps to achieve stationarity), and q (the number of lagged forecast errors). This model is particularly effective for univariate time series forecasting where patterns like trends or seasonality are present.
All results are reported on a 70/10/20 train/validation/test split for each datasets, except for the ETT datasets which have predefined splits. MSE is reported after all datasets have been scaled by the mean and variance of the training data. Both autoregressive models have only one tunable hyperparameter (number of lags). Similarly, the linear model has only one tunable hyperparameter (number of training epochs).
The baselines as described can handle only univariate time series, while all of the benchmark datasets are multivariate (multiple channels). These baselines are trained under channel independence: each channel of a time series is treated independently. While channel independence fails to take into account cross-channel dependencies, we note that developing methods that leverage cross-channel dependencies for a variable number of channels remains an open problem.
C.3.1 Time-LLM
Due to differences in the results reported for Time-LLM between the original paper and reproductions in several works (Goswami et al., 2024; Ekambaram et al., 2024; Pan et al., 2024; Tan et al., 2024), we to attempt our own partial reproduction in order to determine whether to include their numbers. To do so, we used code provided by the authors and looked at ETTh1 with input lengths from ${96,192}$ and ILI with input lengths from ${24,36,48,90}$ (going beyond this was infeasible due to the resources required to finetune LLaMA-7B).
Our reproduced MSE for ETTh1 with an input length of 96 was 0.405 as compared to that of 0.362 reported by Jin et al. (2024), while for input length of 192 it was 0.431 vs. 0.398. On ILI the average increase in MSE across time horizons from their reported results to our reproduction was more than 0.4. Due to these large discrepancies we chose to exclude their model from our analysis.
C.3.2 Closed-source models
In Table 14 we consider three additional FMs that either do not report complete set of results (Moirai) on all 7 datasets or are closed-source (Toto and LLM4TS). While the performance of Auto-AR is comparably to that of LLM4TS, Toto dominates the three aggregated metrics we considered in Table 15; impressively, it does this despite being zero-shot. However, a look at Figure 4 reveals that Toto is not superior across all six tasks, with the aggregated metrics being strongly boosted by its dramatically better performance on one of them (ETTh2). Thus, while its ZS performance is quite good, it is unclear whether this result would be maintained with additional tasks.
Model Partial setting (6 datasets / 24 tasks)
Model Pretraining Avg. Avg. Median

### v2

@@ 4.3 Time series
Supervised Auto-ARIMA 10 0 0.896 12.82 0.00 0.00
Models AR 513 0 0.556 6.57 31.17 24.31
@@
Table 3: Aggregate forecasting performance across seven tasks, demonstrating that simply tuning a classical AR model is competitive with state-of-the-art FMs while using no pretraining data and tens of thousands of times fewer parameters. For Mean / Median %Imp. we report % improvement over Auto-ARIMA, and for Auto-AR, the model size refers to the largest configuration across tasks.
@@
Our last domain is time series, which has many FMs that use the pretrain-then-fine-tune workflow, such as GPT4TS (OFA) (Zhou et al., 2023a), LLM4TS (Chang et al., 2023), MOMENT (Goswami et al., 2024), TEST (Sun et al., 2023), S2IP-LLM (Pan et al., 2024), CALF (Liu et al., 2024), TTM (Ekambaram et al., 2024), and Time-LLM (Jin et al., 2024), and others that evaluate in a zero-shot (ZS) regime, such as TEMPO (Cao et al., 2024), TimesFM (Das et al., 2024), Moirai (Woo et al., 2024), and Toto (Cohen et al., 2024). We do not include four of these FMs in our main analysis (Table 16): for Time-LLM, this is because their results have been hard-to-reproduce for both us and past efforts (c.f. Appendix C.3.1), while the three others—Moirai, LLM4TS, and Toto—either evaluate on only a subset of our chosen tasks or are closed-source (or both); we list their reported numbers in Table 14. As discussed in Appendix C.3.2, Moirai underperforms Auto-AR while LLM4TS performs roughly on par with TTM (A). The most recent FM, Toto, has strong aggregate metrics, mainly due to a dominant performance on a single task, ETTh2 (c.f. Figure C.3.2). As a result we do not view these excluded FMs as significantly affecting our conclusions. We also did not include another FM-like approach, TOTEM (Talukder et al., 2024), because of its focus on multi-task training rather than pretraining and because its multi-task results underperform its single-task method, which if anything supports our claim about the limited (realized) benefits of transfer learning in time series.
We focus on long-horizon forecasting, which has a standard set of datasets (Goswami et al., 2024, Table 11), of which we consider seven. [^5] The two we do not consider, Exchange and ILI, are not evaluated on by most time series FMs. Each consists of four settings corresponding to different time horizons, so in total this yields twenty-eight tasks. We compute aggregate metrics using RMSE, not MSE, so that performance scales linearly with prediction error; this choice has no effect on average rank. Lastly, note that since our focus is on supervised baselines, our main comparison with zero-shot models (Table 16) is in a less challenging setting than the one they report numbers for. However, AR can also be used in the ZS setting by training on just the history provided for the test series; as reported in Section 5.5, this is competitive with open-source ZS FMs at short time horizons.
Baselines:
@@
As baselines we mainly use linear forecasters, including the classical (untuned) linear auto-regression (AR), the automated workhorse Auto-ARIMA (Hyndman & Khandakar, 2008), the more recent DLinear (Zeng et al., 2023), and our own workflow Auto-AR described in Section 3.2. Lastly, we also evaluate our other approach, DASHA, on six of the tasks (c.f. Table 14).
Results:
@@
Table 16 shows that on the full seven-dataset evaluation our Auto-AR workflow always attains competitive performances across all aggregate metrics considered, and in particular attains the best median improvement over Auto-ARIMA. Specifically, our Auto-AR workflow achieves competitive performances with two other recent time series FMs, namely MOMENT and S2IP-LLM, and outperforms the rest of the FMs. Although TTM surpasses all other methods across three aggregated metrics, the improvements remain relatively marginal. Thus, in this the domain as well we find that increases in pretraining dataset size and model scale have not yet resulted in significant performance improvements. Notably, even untuned AR, which uses no differencing and a large lookback window, is quite effective, doing better than zero-and-few-shot FMs across all metrics.
@@ 5.1 The importance of diverse, well-tuned, and domain-specific baselines
5.2 The importance of benchmarks that address actual FM use-cases
A common argument in favor of foundation models is utility in low-data regimes, where transfer might be necessary to cope with limited supervision. However, to demonstrate usefulness via this argument, FMs must be evaluated in data-scarce regimes rather than the usual full-data settings they (and thus also we) generally consider. Furthermore, as demonstrated in Figure 5.2, our supervised baselines remain competitive even when only 20% of the data is provided to each task. Together with the zero-shot Auto-AR results (c.f. Section 5.5), this indicates that significant effort needs to be put in to designing benchmarks that elucidate the claimed potential of existing FMs.
Figure 3: Performance of the best FMs when given only 20% as much fine-tuning data (c.f. Table 19). Supervised baselines are competitive even in this data-scarce regime: in satellite imaging and time series they are outperformed by just one FM, while in genomics they are beaten by two. Notably the worst of the three genomics FMs in the full setting does best with less data.
Figure 4: Tuning costs (excluding FM pretraining but including HPO for all methods and NAS for DASHA) and downstream fine-tuning / re-training costs (c.f. Table 18). Apart from Caduceus, supervised models are much cheaper, and for time series the entire Auto-AR pipeline is 3.5×3.5$\times$3.5 × faster than fine-tuning any FM once (n.b. tuning costs for time series FMs are unknown).
@@ 5.3 Computational efficiency considerations
While computational efficiency is not our main focus, we nevertheless note that any performance gains from FMs must also be balanced against any extra cost. In addition to the usually massive costs of pretraining, the resulting models are often very big and costly to use. Indeed, apart from the special case of HyenaDNA and Caduceus, the CNN architectures discovered and trained using DASHA are typically over ten times smaller than FMs in the case of genomics and three to ten times smaller in the case of satellite imaging. Moreover, for time series our Auto-AR approach is quick-to-train and yields simple models with less than 1K parameters—over 2000$\times$ smaller than any FM—while attaining performance that is often competitive even with closed-source models. We visualize the computational advantages of smaller supervised models in Figure 4, where we compare them against the top FMs in each domain. For genomics, fine-tuning most FMs takes roughly the same time as training our baselines, while for satellite imaging, FMs take roughly twice as long. Lastly, in time series our Auto-AR pipeline is over 30$\times$ faster than large FMs and over three times faster than smaller models like TTM. This demonstrates efficiency of the supervised approaches sets a high performance bar that FMs need to clear before they can be deemed useful.
@@ 5.4 The power of tuning kernel sizes and dilation rates
Our results for genomics and satellite imaging are driven by the DASHA workflow, whose crucial component is the tuning of kernel sizes and dilation rate in CNN backbones such as wide ResNet. Its success demonstrates that the procedure is an effective surrogate for human-driven model development, enabling the automated discovery of the types of diverse, domain-specific baselines stressed in Section 5.1. To understand this further, we study whether the architecture search component selects different kernel sizes and dilation rates for different tasks, and whether it does so in a consistent manner. Specifically, we run DASHA on three of the smaller datasets in the NT benchmark with fifteen different random seeds, construct eighteen-dimensional vectors of the discovered kernel sizes and dilation rates assigned to each of the nine layers, and project these to two dimensions using principal component analysis (PCA). The result in Figure 5.4 reveals that the architectures are clustered by task, demonstrating that the procedure selects different but consistent-within-task kernel parameters. This visualization suggests that architecture search is a useful surrogate for model development, and consequently that the DASHA workflow may also be useful for automating similar studies and baselining FMs in other domains with high-dimensional, unstructured data.
Figure 5: PCA visualization of architectures discovered for three different tasks when DASHA is run multiples times. Clustering across tasks reveals the within-task consistency of the architecture search component and the utility of diverse models as baselines.
Figure 6: Zero-shot performance on the subset of time series tasks with Moirai and Toto evaluations. At shorter horizons, Auto-AR trained on only the history given by the test series (and thus also given to ZS FMs) is competitive with all FMs except the closed-source Toto.
@@ 5.5 The surprising effectiveness of linear auto-regression
As described in Appendix F, Auto-AR also turns out to be easily extended to and competitive in the zero-shot setting, enabling a direct comparison with ZS FMs (c.f. Figure 6). The method involves training a short-lookback AR model on subsequences of the history of the length 512 provided in the ZS regime. This establishes what we believe should be the default supervised baseline for time series forecasting in the ZS setting. We provide full evaluations in Tables 2021.
@@ 5.6 Limitations
While our findings are significant according to measures set by past work, they should not be misinterpreted to address all possible scenarios where FMs may be useful; for example, genomics FMs are often used for exploratory science and not prediction. We are also of course computationally limited and there are many other domains where FMs have been pretrained, and even in our three there are other tasks beyond classification and forecasting. Nevertheless, our evaluation is extensive—over fifty tasks and thirty FMs—and so the results provide at least a strongly indicator w.r.t. the state of a field that uses benchmark performance to motivate and justify pretraining.
@@ Acknowledgments
We thank Mononito Goswami, Esther Rolf, and Stephan Xie for useful feedback. This work was supported in part by NSF grants IIS1705121, IIS1838017, IIS2046613, IIS2112471, a TCS Presidential Fellowship, and funding from Meta, Morgan Stanley, Amazon, Google, and Scribe. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of any of these funding agencies.
@@ C.1 Genomics
We include all FMs listed in Dalla-Torre et al. (2023, Supplementary Table 6) with addition of the two recently released models, Caduceus (Schiff et al., 2024) and Gena-LM (Fishman et al., 2024); note that the last row of tasks in the NT paper(promoter and splice sites) are mislabeled, but we infer an order in combination with previous information obtained from a (since-deleted) Huggingface leaderboard. [^6] See https://huggingface.co/spaces/InstaDeepAI/nucleotide_transformer_benchmark. In alignment with the leaderboard, we apply a 0.1 validation split for DASHA during our evaluation. Additionally, we use an architecture set that includes both Wide ResNet and UNet for the search with DASHA on these datasets. We use batch size$=$ 128 for all datasets, and cross entropy loss for all the training and finetuning. Individual scores for each task in the benchmark are provided in Tables 9 and 8.
@@
Table 8: Regulatory Elements and RNA Production tasks. “-” indicates unknown quantities.
@@
Table 9: Histone Modification tasks
@@
Table 11: Regulatory Elements and RNA Production for GENA-LM. (all scores are obtained using MCC). The last column shows the average MCC across all 18 datasets in Tables 10 and 11.
@@ C.2 Satellite imaging
Training on satellite datasets requires relatively large computational resources due to the high number of channels and the size of the datasets. To ensure a fair comparison, we fine-tuned all the foundation models ourselves by sweeping across a fixed set of base learning rates $[5e-3,2e-3,2e-4,4e-5]$. We then calculate the actual learning rate from base learning rate following previous work by $lr=base\_lr\cdot\frac{batch\_size}{256}$. This approach ensures that approximately the same amount of resources were used as during the DASHA tuning process, allowing for a balanced evaluation of model performance. For training and finetuning, we universally use batch_size = 16 and loss function as cross entropy with 0.1 label smoothing for single label classification, and multi-label soft margin loss for multilabel classification. Individual scores for each task are provided in Table 12.
@@
It is also important to note that SatMAE only accepts 3-channel and 12-channel inputs, while CROMA is limited to 12-channel inputs. GeoBench, however, includes a wide range of tasks with varying numbers of input channels, ranging from 3 to 18. Despite these differences, we include all datasets in our evaluation because they are valuable benchmarks in the satellite image domain, and it is crucial for FMs in this field to generalize across diverse datasets. For datasets where the input size does not match the model requirements, we pad missing channels with zeros and prune any extra channels. However, to ensure a fair comparison, in addition to reporting the average scores across all datasets, we also provide average scores excluding m-pv4ger and m-forestnet, where missing channels may affect the performance of the FMs. The aggregate scores excluding m-pv4ger and m-forestnet are presented in Table 13.
Model Average Average Mean Median
Score $\uparrow$ Rank $\downarrow$ %Imp.$\uparrow$ %Imp.$\uparrow$
SatMAE-Base 77.71 6.00 6.74 4.56
SatMAE-Large 78.44 4.07 7.87 6.75
Foundation GFM 76.95 5.57 5.40 4.08
Models SwinT-Base 76.12 6.50 3.98 1.43
CROMA-Base 77.98 3.57 6.95 5.30
CROMA-Large 79.06 2.14 8.84 6.26
ResNet50 73.25 9.00 -0.27 -0.38
Supervised Wide ResNet 73.90 8.00 0.00 0.00
Models UNet 75.29 6.57 2.33 0.87
DASHA (our workflow) 78.00 3.57 7.02 5.16
@@ C.3 Time series
Long horizon forecasting for time series is the following task: at every timestep $t$, take the $L$ historical observations at times $(t-L+1,...,t)$ for each channel and predict the next $H$ observations $(t+1,...,t+H)$ for every channel. Following the literature, we evaluate on $H\in\{24,36,48,60\}$ on ILI and $\{96,192,336,720\}$ otherwise. For most methods, we report results for $L=512$. All results are reported on a 70/10/20 train/validation/test split for each datasets, except for the ETT datasets which have predefined splits. MSE is reported after all datasets have been scaled by the mean and variance of the training data.
In addition to the results for DASHA, Auto-AR, DLinear, and the FMs, we evaluate the performance of two simple baselines:
  1. 1.
    AR (d=0): The vanilla autoregressive model (Box & Jenkins, 1976) predicts the (scalar) value of a time series at $t+1$ as a linear combination of the last $L$ timesteps and a constant, i.e. $\hat{x}_{t+1}=\alpha_{0}+\alpha_{1}x_{t}+\alpha_{2}x_{t-1}+...+\alpha_{L}x_{t-L+1}$ for learnable parameters $\alpha_{0},...,\alpha_{L}$. We fit these parameters using standard maximum likelihood techniques. Here “d=0” denotes no differencing.
  2. 2.
    Auto-ARIMA: ARIMA is a statistical method used for time series forecasting that combines three components: AutoRegressive (AR), Integrated (I), and Moving Average (MA). The AR component models the relationship between an observation and its lagged (past) values, assuming that past values have a linear influence on future ones. The Integrated component applies differencing to the data to remove trends or seasonality, making the time series stationary by stabilizing its mean over time. The MA component models the relationship between an observation and the residual errors from a moving average model applied to previous observations. ARIMA is characterized by three parameters: p (the number of lag observations), d (the number of differencing steps to achieve stationarity), and q (the number of lagged forecast errors). This model is particularly effective for univariate time series forecasting where patterns like trends or seasonality are present. To tune these parameters we use the popular Auto-ARIMA approach of Hyndman & Khandakar (2008) discussed in Section 3.2.
As described, all of our baselines can handle only univariate time series, while all of the benchmark datasets are multivariate (multiple channels). These baselines are trained under channel independence: each channel of a time series is treated independently. While channel independence fails to take into account cross-channel dependencies, we note that developing methods that leverage cross-channel dependencies for a variable number of channels remains an open problem.
@@
TEMPO (Zero Shot) 0.211 0.254 0.292 0.37 0.178 0.198 0.209 0.279 - - - - 0.476 0.496 0.503 0.538
TimesFM (Zero Shot) 0.122 0.169 0.242 0.391 0.119 0.137 0.158 0.206 - - - - 0.327 0.353 0.378 0.42
@@
Table 14: Time series forecasting tasks (MSE). “-” indicates unknown quantities.
Model ETTh1 ETTh2 ETTm1 ETTm2
Horizon 96 192 336 720 96 192 336 720 96 192 336 720 96 192 336 720
GPT4TS (OFA) 0.397 0.418 0.433 0.456 0.342 0.389 0.407 0.441 0.346 0.372 0.394 0.421 0.262 0.301 0.341 0.401
TEST (Few shot) 0.457 0.519 0.531 0.594 0.374 0.433 0.44 0.48 0.401 0.426 0.444 0.501 0.262 0.302 0.341 0.419
LLM4TS 0.394 0.412 0.422 0.444 0.332 0.377 0.396 0.425 0.343 0.366 0.385 0.419 0.254 0.292 0.326 0.38
MOMENT 0.41 0.426 0.437 0.472 0.345 0.386 0.408 0.439 0.349 0.368 0.384 0.416 0.26 0.297 0.328 0.387
TTM (B) - - - - - - - - - - - - - - - -
TTM (A) - - - - - - - - - - - - - - - -
$S^{2}$-LLM 0.396 0.42 0.431 0.458 0.34 0.385 0.406 0.436 0.346 0.365 0.39 0.418 0.257 0.299 0.33 0.39
CALF 0.389 0.423 0.436 0.467 0.331 0.38 0.394 0.426 0.349 0.375 0.399 0.438 0.256 0.297 0.339 0.393
TEMPO (Zero Shot) 0.406 0.421 0.43 0.451 0.353 0.389 0.408 0.44 0.424 0.432 0.467 0.509 0.267 0.304 0.345 0.395
Moirai (Zero Shot) 0.402 0.419 0.429 0.444 0.334 0.373 0.393 0.416 0.383 0.402 0.416 0.437 0.282 0.318 0.355 0.41
Toto (Zero Shot) 0.366 0.368 0.399 0.424 0.197 0.231 0.260 0.355 0.328 0.353 0.389 0.429 0.27 0.315 0.319 0.374
DLinear 0.399 0.416 0.443 0.49 0.353 0.418 0.465 0.551 0.343 0.365 0.386 0.421 0.26 0.303 0.342 0.421
Auto-ARIMA 0.537 0.566 0.587 0.608 0.368 0.418 0.454 0.467 0.659 0.683 0.701 0.725 0.296 0.34 0.381 0.438
Auto-AR 0.388 0.41 0.423 0.448 0.334 0.376 0.402 0.437 0.343 0.364 0.384 0.418 0.251 0.29 0.325 0.381
Model Weather Electricity ILI Traffic
Horizon 96 192 336 720 96 192 336 720 96 192 336 720 96 192 336 720
GPT4TS (OFA) 0.212 0.248 0.286 0.337 0.238 0.251 0.266 0.297 0.881 0.892 0.884 0.957 0.282 0.29 0.294 0.312
LLM4TS 0.196 0.238 0.277 0.329 0.223 0.24 0.258 0.292 - - - - 0.259 0.265 0.275 0.292
TEST (Few shot) 0.213 0.263 0.282 0.328 0.235 0.255 0.275 0.311 - - - - 0.317 0.3 0.31 0.338
MOMENT 0.209 0.248 0.285 0.336 0.233 0.247 0.264 0.295 1.114 1.092 1.098 1.126 0.282 0.287 0.292 0.31
TTM (B) - - - - - - - - - - - - - - - -
TTM (A) - - - - - - - - - - - - - - - -
$S^{2}$-LLM 0.195 0.235 0.28 0.326 0.23 0.247 0.266 0.287 - - - - 0.274 0.282 0.289 0.301
CALF 0.204 0.25 0.291 0.352 0.238 0.252 0.267 0.303 - - - - 0.268 0.278 0.281 0.3
TEMPO (Zero Shot) 0.254 0.298 0.332 0.379 0.276 0.293 0.309 0.355 - - - - 0.343 0.355 0.356 0.376
TimesFM (Zero Shot) 0.122 0.169 0.242 0.391 0.119 0.137 0.158 0.206 - - - - 0.327 0.353 0.378 0.42
Moirai (Zero Shot) 0.212 0.25 0.282 0.322 0.299 0.31 0.323 0.347 - - - - - - - -
Toto (Zero Shot) 0.223 0.267 0.291 0.356 0.212 0.232 0.249 0.291 - - - - - - - -
DLinear 0.237 0.282 0.319 0.362 0.237 0.249 0.267 0.301 1.081 0.963 1.024 1.096 0.282 0.287 0.296 0.315
Auto-ARIMA 0.267 0.341 0.426 0.53 0.441 0.468 0.485 0.54 - - - - 0.753 0.819 0.932 1.243
Auto-AR 0.223 0.26 0.294 0.341 0.233 0.247 0.265 0.301 0.944 0.948 0.947 0.958 0.28 0.286 0.294 0.319
Table 15: Time series forecasting tasks (MAE). “-” indicates unknown quantities.
Model Average Average Mean Median
MAE $\downarrow$ Rank $\downarrow$ %Imp.$\uparrow$ %Imp.$\uparrow$
GPT4TS 0.337 4.54 33.01 30.07
Foundation TEST (Few Shot) 0.370 6.36 27.35 26.89
Models MOMENT 0.336 3.68 33.32 30.19
S2IP-LLM 0.331 2.54 34.22 32.68
CALF 0.335 3.68 33.50 29.63
Supervised DLinear 0.350 5.61 29.87 26.10
Models Auto-ARIMA 0.553 7.82 0.00 0.00
Auto-AR (our workflow) 0.333 2.71 33.86 29.46
Table 16: Aggregate performance on time series tasks across seven tasks based on MAE.
Model Pretraining Average Mean Median